### v1

# Forget What You Know about LLMs Evaluations - LLMs are Like a Chameleon

Abstract
Large language models (LLMs) often appear to excel on public benchmarks, but these high scores may mask an overreliance on dataset-specific surface cues rather than true language understanding. We introduce the Chameleon Benchmark Overfit Detector (C-BOD), a meta-evaluation framework that systematically distorts benchmark prompts via a parametric transformation and detects overfitting of LLMs. By rephrasing inputs while preserving their semantic content and labels, C-BOD exposes whether a model’s performance is driven by memorized patterns. Evaluated on the MMLU benchmark using 26 leading LLMs, our method reveals an average performance degradation of 2.15% under modest perturbations, with 20 out of 26 models exhibiting statistically significant differences. Notably, models with higher baseline accuracy exhibit larger performance differences under perturbation, and larger LLMs tend to be more sensitive to rephrasings indicating that both cases may overrely on fixed prompt patterns. In contrast, the Llama family and models with lower baseline accuracy show insignificant degradation, suggesting reduced dependency on superficial cues. Moreover, C-BOD’s dataset- and model-agnostic design allows easy integration into training pipelines to promote more robust language understanding. Our findings challenge the community to look beyond leaderboard scores and prioritize resilience and generalization in LLM evaluation.
Forget What You Know about LLMs Evaluations - LLMs are Like a Chameleon
Nurit Cohen-Inger1, Yehonatan Elisha2, Bracha Shapira1, Lior Rokach1, Seffi Cohen1 1Ben Gurion University  2Tel Aviv University
1 Introduction
Large Language Models (LLMs) have achieved impressive results on a wide range of NLP tasks Chang et al. (2024). Consequently, hundreds of benchmarks have been established to track progress and evaluate model capabilities Lu et al. (2024); Liang et al. (2022). However, the rapid proliferation of LLMs and the frequent use of public leaderboards raise concerns about the robustness of these evaluation practices Castillo-Bolado et al. (2024). Specifically, as benchmark data becomes more widely recognized, models may learn to exploit surface patterns or spurious correlations, rather than exhibit genuine language understanding. This issue can lead to deceptively high scores that do not reflect true progress.
In this paper, we examine whether LLMs rely excessively on benchmark-specific cues potentially overfitting to the patterns inherent in widely published evaluation benchmarks and explore systematic methods to detect and mitigate this behavior. In other words, are LLMs prone to overfitting on popular benchmarks, and what underlying factors contribute to this phenomenon?
To answer this question, we introduce Chameleon Benchmark Overfit Detector (C-BOD), a framework that reveals how heavily a model depends on the exact wording or structure of a test set. By introducing controlled textual distortions to benchmark prompts at varying intensities (defined by a distortion parameter $\mu$), as demonstrated in Figure 1, our method reveals whether high performance is built on superficial patterns. Notably, our framework requires only the evaluation set, without accessing the model’s training data or architecture. Unlike conventional leaderboards that solely track performance, our meta-evaluation framework acts as a safeguard ensuring that high scores do not stem from superficial memorization of benchmark cues.
Figure 1: An example demonstrating the C-BOD method. The original question (top) is perturbed (bottom) while preserving the semantic meaning and correct answer options. The model correctly answers the original question but fails on the perturbed version, suggesting potential overfitting. Changes in the perturbed question are highlighted in bold.
Our Contributions:
  1. 1.
    Robust Overfitting Detection with Statistical Significance. Our framework computes the performance difference $\Delta_{\mu}$ between original and perturbed prompts and confirms its statistical significance, ensuring that observed differences indeed indicate overfitting rather than chance variations.
  2. 2.
    New Findings For LLM Community Our extensive analysis reveals new trends regarding how LLMs function with respect to their number of parameters and baseline performance.
  3. 3.
    Extensive Empirical Validation. We apply our method to multiple LLM families of various architectures and parameter sizes. Even modest textual distortions cause significant performance differences in most models, providing strong empirical evidence that overfitting is widespread.
  4. 4.
    Publicly Available Benchmarks and Code. We release rephrased versions of the widely used MMLU evaluation set under different distortion levels ($\mu$). These resources enable the community to adopt more robust, surface-invariant tests for reliable LLM assessment with our method reproducible code.
  5. 5.
    Blueprint for Iterative Overfit Mitigation. Beyond detection, these $\mu$-based rephrasings can be integrated into model training or fine-tuning pipelines. Regularly exposing models to diverse prompt variations helps reduce reliance on benchmark-specific phrasing, thus promoting more generalizable language understanding.
2 Related Work
2.1 Benchmark Evaluation and Overfitting
LLMs have achieved impressive results on many benchmarks. This success has driven the development of comprehensive evaluation suites such as BIG-Bench Srivastava et al. (2022) and HELM Liang et al. (2022). MMLU benchmark set Hendrycks et al. (2020) evaluates question answering across 57 subjects—including STEM, humanities, and social sciences, while Zhang et al. (2024a) introduced 25 enterprise-focused datasets covering domains like finance, legal, cybersecurity, and climate sustainability for tasks such as classification, NER, and summarization. Another recent resource, JUDGE-BENCH Bavaresco et al. (2024), comprises 20 NLP datasets that assess models against human judgments. We focus on MMLU because of its widespread adoption and comprehensive domain coverage Wang et al. (2024), which makes it particularly effective for exposing overfitting to canonical prompt structures.
While these benchmarks have been critical for comparing new models’ versions, recent studies warn that publicly released evaluation sets can become less reliable over time due to overexposure and memorization Yu et al. (2024); Chang et al. (2024). In some cases, LLMs learn superficial patterns specific to well-known datasets, boosting performance without reflecting genuine semantic or conceptual understanding. Kiela et al. (2021) further emphasize the need for continuously refreshing benchmarks to ensure real progress in language understanding. For example, OpenAI’s GPT models have shown steady improvement on MMLU: GPT-3 achieved approximately 43% accuracy in 2020 Brown et al. (2020), rising to nearly 70% with GPT-3.5 in 2022111https://cdn.openai.com/papers/GPT-4-Technical-Report.pdf, and reaching 86% with GPT-4 in 2023  [^2] https://cdn.openai.com/papers/GPT-4-Technical-Report.pdf.
Memorization in LLMs has been widely studied Kiyomaru et al. (2024); Biderman et al. (2024), with larger models especially prone to retaining training data verbatim Carlini et al. (2022). This phenomenon can inflate performance metrics while obscuring genuine model capabilities. Moreover, several works highlight training-set contamination, where test samples appear exactly or as near-duplicates in the training data, as another crucial form of overfitting Deng et al. (2023); Yao et al. (2024), leading to overly optimistic performance estimates Yang et al. (2023).
2.2 Gap in Current Work
Researchers have introduced various methods to detect or mitigate training contamination: Finding the N-gram overlap (e.g., 13-grams or 50-character matches) between training and test data Brown et al. (2020); OpenAI (2023), though it can miss semantically equivalent rephrasing. Embedding similarity search Reimers (2019) that uses transformer-based embeddings to identify semantically close training-test pairs Lee et al. (2023). Decoding Matching probes the model by providing partial test prompts and measuring how likely it is to complete them exactly Li (2023) or completing missing words Deng et al. (2023). A recent study presented an overfit detection of editing knowledge to a LLM Zhang et al. (2025).
Although these studies have focused on detecting training data contamination or focusing on additional knowledge, they lack with addressing a critical issue: overfitting to benchmark-specific artifacts. In many cases, LLMs may never see the test data during training yet still learn to rely on superficial cues unique to a benchmark’s canonical format. Existing techniques such as n-gram overlap and embedding similarity fail to capture this subtle form of overfitting. In contrast, our approach explicitly quantifies the dependence of a model’s performance on the precise phrasing and structure of evaluation prompts, thereby filling this gap in current evaluation methodologies. By systematically applying a controllable distortion parameter to evaluation prompts, without requiring additional training or access to training data, our method shows how performance metrics degrade under textual perturbations, providing a robust means of diagnosing and mitigating broader overfitting behavior.
3 Method
Let $\mathcal{D}$ denote a benchmark dataset with N samples, and $\mathcal{E}$ a LLM to be evaluated with respect to a given performance function $\mathcal{M}$. Our goal is to detect whether $\mathcal{E}$ exhibits overfitting to $\mathcal{D}$. Figure 2 provides an overview of our proposed method, Chameleon Benchmark Overfit Detector (C-BOD). C-BOD employs a rephrasing transformation to generate a perturbed dataset from $\mathcal{D}$, evaluates on both the original and perturbed datasets, and applies a statistical test to assess whether performance discrepancies indicate overfitting. The following subsections detail each component of C-BOD.
Figure 2: High-level pipeline of our parametric approach. The original dataset $\mathcal{D}$ is passed through the distortion operator $T_{\mu}$ to form $\mathcal{D}_{\mu}$. Both sets are evaluated by a LLM, and differences in performance are used to quantify overfitting.
3.1 C-BOD rephrased dataset generation
To systematically introduce textual variations, C-BOD utilizes a rephrasing tool, denoted as ${T}$, which uses as a distortion operator to generate a perturbed dataset $\mathcal{D}_{\mu}$ from $\mathcal{D}$. This operator is parameterized by $\mu$ (temperature), which controls the extent of textual modification, ranging from low (e.g., 0.1 for minimal changes like synonym substitution) to moderate (e.g., 1.0 for rewording and sentence fragment reordering) and high (e.g., 1.5 for aggressive modifications such as question reformulation). We define:
$T_{\mu}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$
Given a prompt $x_{i}$, the distortion operator produces a perturbed prompt $x_{i}^{\prime}=T_{\mu}(x_{i})$. The perturbed dataset is then constructed as:
$\mathcal{D}_{\mu}\;=\;\left\{\,\bigl{(}x_{i}^{\prime},\,y_{i}\bigr{)}\;\middle|\;(x_{i},y_{i})\in\mathcal{D}\right\}$
Although each pair $(x_{i}^{\prime},y_{i})$ in the perutbed dataset remains semantically equivalent to $(x_{i},y_{i})$ in the original dataset, the textual variations introduced by $T_{\mu}$ can disrupt purely memorized mappings from surface patterns to correct labels. This step presented in Lines 5-6 of Algorithm 1.
3.2 Evaluating the Impact of Distortion
To assess the impact of distortion, we evaluate $\mathcal{E}$ using a performance function, $\mathcal{M}$. This function evaluates $\mathcal{E}$ based on a given ground truth or prompt $y_{i}$, considering two versions of an input: the original $x_{i}\in\mathcal{D}$ and the perturbed version $x_{i}^{\prime}\in\mathcal{D}_{\mu}$, where $i$ denotes the index of a sample in the dataset. Specifically, $\mathcal{M}$ is a boolean function that takes as input the model $\mathcal{E}$ and two data pairs, $(x_{i},y_{i})$ and $(x_{i}^{\prime},y_{i})$, and returns whether the model performs better on the original input than on the perturbed one. The function is formulated as follows:
$\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x^{\prime}_{i},y_{i})=\begin{cases}1,&\text{if }P(\mathcal{E},x_{i},y_{i})\\ &\quad>P(\mathcal{E},(x^{\prime}_{i},y_{i}),\\ 0,&\text{otherwise.}\end{cases}$
where $P(\mathcal{E},x,y)$ represents the performance score of model $\mathcal{E}$ on input $x$ with reference to ground truth $y$. This formulation is designed to be generalizable across different evaluation metrics and natural language understanding (NLU) tasks. The performance difference between the original set and the perturbed set is then calculated as:
$\Delta_{\mu},b=\sum_{i=0}^{N}\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x_{i}^{\prime},y_{i}))$ (1)
The performance difference between the perturbed set and the original set is then calculated as:
$c=\sum_{i=0}^{N}\mathcal{M}(\mathcal{E},(x_{i}^{\prime},y_{i}),(x_{i},y_{i}))$ (2)
A large positive $\Delta_{\mu}$ indicates a significant performance decline due to textual perturbations, suggesting that $\mathcal{E}$ may be overly reliant on surface-level patterns rather than exhibiting robust generalization. Notably, this approach remains metric-agnostic, making it applicable to a wide range of evaluation measures. This step presented in Lines 7-8 of Algorithm 1.
3.3 Statistical Validation
To assess the statistical significance of performance differences, we employ McNemar’s test McNemar (1947), which is specifically designed for paired data. This test evaluates whether the discrepancies between two related sets of classification outcomes, correct and incorrect predictions, are significant. In our context, McNemar’s test is well-suited for comparing each pair of samples $(x_{i},y_{i})\in D$ and $(x_{i}^{\prime},y_{i})\in D_{\mu}$, we record whether $\mathcal{E}$ classifies them correctly and aggregate into b (original is better) and c (perturbed is better) as presented in Equation 1, Equation 2. The McNemar statistic is then calculated as:
$\chi^{2}=\frac{(b-c)^{2}}{b+c}$ (3)
We derive a $p$-value from the chi-squared distribution (with df=1, i.e., one degree of freedom), rejecting the null hypothesis if $p<\alpha$. A significant result with $b>c$ indicates a genuine performance difference due to the transformation, suggesting evidence of overfitting. This step presented in Lines 10-19 of Algorithm 1.
Algorithm 1 Chameleon Benchmark Overfit Detector
1:
2:$\mathcal{D}$: Original benchmark dataset of size $N$,
$\mathcal{E}$: LLM,
4:$\mu$: Distortion parameter,
5:$T_{\mu}$: Transformation operator,
$\mathcal{M}$: Performance function (returns 1 if the first input is better, 0 otherwise),
7:$\alpha$: Significance level.
8:C-BOD Computation:
9:$b,c\leftarrow 0$
10:$D_{\mu}\leftarrow\{\}$
11:for each $x_{i}\in\mathcal{D}$ do
12:     $x^{\prime}_{i}\leftarrow T_{\mu}(x_{i})$
13:     $D_{\mu}\leftarrow D_{\mu}\cup\,x_{i}^{\prime}$
14:     $b\leftarrow b+\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x_{i}^{\prime},y_{i}))$
15:     $c\leftarrow c+\mathcal{M}(\mathcal{E},(x_{i}^{\prime},y_{i}),(x_{i},y_{i}))$
16:end for
17:$\chi^{2}\leftarrow\dfrac{(b-c)^{2}}{b+c}$
18:$p\leftarrow\text{p-value}(\chi^{2},\text{df}=1)$
19:if $p<\alpha$ then
20:     if $b>c$ then
21:         $Overfit\_Flag\leftarrow True$
22:     else
23:         $Overfit\_Flag\leftarrow False$
24:     end if
25:else
26:     $Overfit\_Flag\leftarrow False$
27:end if
28:return $Overfit\_Flag$
4 Experimental Setting
In this section, we describe the experimental setup used to evaluate our overfitting detection framework. We detail the benchmark dataset, the procedure for generating perturbed inputs, the LLMs under evaluation, implementation specifics, and the evaluation metrics.
4.1 Dataset and Rephrasing Process
Our experiments use the MMLU benchmark Hendrycks et al. (2020), which comprises multiple-choice questions spanning 57 subjects. The broad coverage and public availability of MMLU make it an ideal candidate for assessing general knowledge and the degree to which LLMs overfit canonical prompt formats. The MMLU dataset is distributed under the MIT License, which allows for free use, modification, and distribution as long as the original copyright notice and license terms are maintained. We generate a perturbed versions of the original dataset to probe overfitting, following the methodology described in Section 3. We used DeepSeek 3 to create the transformed version of each question. We generate the perturbed dataset using $\mu=1.0$ (the default temperature parameter).
These perturbations include synonym substitutions, sentence reordering, and the insertion of distractor phrases, while preserving the original semantic meaning and correct answers. Automated formatting checks and manual audits (performed on approximately 10% of the samples) ensure that the integrity of the questions is maintained. For example, an original question:“The coronal suture joins the?” is rephrased as:“Which bones does the coronal suture connect?”. The perturbed dataset, denoted by $\mathcal{D}_{\mu}$, is released alongside our code for reproducibility  [^3] https://github.com/SeffiCohen/CBOD.
4.2 Models Under Evaluation
Table 1 provides an overview of the LLMs evaluated in our experiments. Our study covers a diverse set of architectures and parameter scales ranging from 1B to 236B parameters. This broad selection enables an in-depth analysis of how both architectural choices and model scale affect robustness to prompt perturbations.
Table 1: Overview of the evaluated LLMs. Models are grouped by family, model version, and the number of parameters (in billions).
Family Version Params
Qwen Qwen2.5 1.5B Yang et al. (2024) 1.5
Qwen2.5 3B 3
Qwen2.5 7B 7
Qwen2.5 32B 32
Qwen2.5 72B 72
Llama 3 Llama 3.2 1B Dubey et al. (2024) 1
Llama 3.2 3B 3
Llama 3.1 8B 8
Gemma Gemma 2 2B Team et al. (2024) 2
Gemma 7B 7
Gemma 27B 27
Phi Phi 3.5 4B Abdin et al. (2024) 4
Phi 4 15B 15
DeepSeek DeepSeek 7B Bi et al. (2024) 7
DeepSeek V2 16B 16
DeepSeek 236B 236
Yi Yi 6B Young et al. (2024) 6
Yi 9B 9
Others Apollo2 7B Zhu et al. (2024b) 7
Aquila 7B Zhang et al. (2024b) 7
Bloomz 7B Zhu (2023) 7
Falcon 7B Almazrouei et al. (2023) 7
Starling 7B Zhu et al. (2024a) 7
Jetmoe 8B Shen et al. (2024) 8
GLM 4 9B GLM et al. (2024) 9
Mistral 8B Jiang et al. (2023) 8
4.3 Implementation Details
All experiments were executed under standardized conditions to ensure reproducibility and fair comparisons:
  1. (1)
    Inference Environment: Most models were accessed via the HuggingFace transformers library using RTX 6000 GPU. DeepSeek 236B model was evaluated using the official API.
  2. (2)
    Dataset Rephrasing Prompt: We instruct the rephrasing tool using the following prompt to generate an alternative version of each question while preserving its original meaning and correct answer: “Rephrase the following question without changing its context or the correct answer: {question}”
  3. (3)
    Query Prompt: For every query, we construct a standardized input by prepending a fixed instruction to the original MMLU question. Importantly, the multiple-choice options remain identical between the original and the rephrased forms. The fixed instruction is:
    “Select the best answer from the given options. Respond with only the letter corresponding to the correct choice.
    Question: {question}”
4.4 Evaluation Metrics
We assess model performance by comparing the original dataset, $\mathcal{D}$, with its perturbed counterpart, $\mathcal{D}_{1.0}$, using the following metrics:
Correct Predictions and Accuracy: For each dataset, we report the number of correct answers and the corresponding accuracy, defined as
Accuracy=#Correct Predictions#Total $\text{Accuracy}=\frac{\#\text{Correct Predictions}}{\#\text{Total Samples}}.$
Absolute and Percentage Performance Difference: The absolute difference in the number of correct answers between $\mathcal{D}$ and $\mathcal{D}_{1.0}$ is denoted by $\Delta_{1.0}$; we also report the relative difference. Statistical Significance: McNemar’s test is applied on the paired predictions to determine whether the performance gap is statistically significant ($p<0.05$)
4.5 Reproducibility
C-BOD source code and datasets, including scripts for data pre-processing, perturbation generation, perturbed datasets, model evaluation, and statistical analysis, are publicly available [^4] https://github.com/SeffiCohen/CBOD. This ensures that our experiments can be independently replicated and verified.
5 Results
5.1 Overall Performance
As shown in Table 2, most models (20 out of 26) exhibit a noticeable drop in performance on the rephrased test set compared to the original, reinforcing our motivation that these LLMs overfit to the standard MMLU format. Notably, the Llama 1B, Llama 3B models maintain relatively stable accuracy, suggesting they are less susceptible to overfitting. We also observed that Falcon 7B, DeepSeek 7B, Qwen 2.5 3B and Jetmoe 8B show statistically insignificant differences, likely due to their lower baseline accuracy. McNemar’s test confirms that the performance declines observed in most models are statistically significant. Notably, no model shows a significant improvement when inputs are rephrased. This indicates that the C-BOD method reliably uncovers model vulnerabilities rather than occasionally yielding unintentional performance gains.
Across all evaluated models, the average drop in accuracy was 2.15%, and when considering only the models with statistically significant differences, this drop increased to 2.72%.
Table 2: Comparison of LLM performance on the original and perturbed MMLU datasets ($\mu=1.0$). The table shows the number of correct answers on each dataset, accuracy, the absolute and percentage performance difference ($\Delta_{1.0}$), statistical significance ($p<0.05$), and whether the model performed better on the original or perturbed dataset. Models are sorted by parameter count (ascending).
Model Name Model Family Par (B) $\mathcal{D}$ Correct $\mathcal{D}$ Accuracy $\mathcal{D}_{1.0}$ Correct $\mathcal{D}_{1.0}$ Accuracy # $\Delta_{1.0}$ $\Delta_{1.0}$ Sig. Better Perf.
Llama 3.2 1B Llama 1 3799 26.98 3802 27.00 -3 -0.08 No Not Sig
Gemma 2 2B Gemma 2 6656 47.28 6552 46.54 104 1.56 Yes Original
Qwen 2.5 3B Qwen 3 5836 41.45 5739 40.76 97 1.66 No Not Sig
Llama 3.2 3B Llama 3 7940 56.40 7989 56.74 -49 -0.62 No Not Sig
Phi 3.5 4B Phi 4 9547 67.81 9325 66.23 222 2.33 Yes Original
Qwen 2.5 1.5B Qwen 5 5137 36.49 4986 35.41 151 2.94 Yes Original
Yi 1.5 6B Yi 6 8899 63.21 8525 60.55 374 4.20 Yes Original
Qwen 2.5 7B Qwen 7 6990 49.65 6810 48.37 180 2.58 Yes Original
Gemma 7B Gemma 7 8173 58.05 8050 57.18 123 1.50 Yes Original
Apollo2 7B Apollo2 7 9547 67.81 9146 64.96 401 4.20 Yes Original
Aquila 7B Aquila 7 4586 32.57 4475 31.78 111 2.42 Yes Original
Bloomz 7B Bloomz 7 6251 44.40 6133 43.56 118 1.89 Yes Original
Falcon 7B Falcon 7 3733 26.51 3718 26.41 15 0.40 No Not Sig
Starling 7B Starling 7 8364 59.41 8179 58.09 185 2.21 Yes Original
DeepSeek 7B DeepSeek 7 6049 42.96 6077 43.16 -28 -0.46 No Not Sig
Llama 3.1 8B Llama 8 6290 44.68 6169 43.82 121 1.92 Yes Original
Mistral 8B Mistral 8 6928 49.21 6691 47.52 237 3.42 Yes Original
Jetmoe 8B Jetmoe 8 6162 43.77 6132 43.55 30 0.49 No Not Sig
GLM 4 9B GLM 9 9674 68.71 9346 66.38 328 3.39 Yes Original
Yi 9B Yi 9 9345 66.38 9180 65.20 165 1.77 Yes Original
Phi 4 15B Phi 15 10776 76.54 10530 74.79 246 2.28 Yes Original
DeepSeek V2 16B DeepSeek 16 7466 53.03 7358 52.26 108 1.45 Yes Original
Gemma 27B Gemma 27 10300 73.16 9903 70.34 397 3.85 Yes Original
Qwen 2.5 32B Qwen 32 11262 80.00 10816 76.82 446 3.96 Yes Original
Qwen 2.5 72B Qwen 72 11456 81.37 11081 78.71 375 3.27 Yes Original
DeepSeek 236B DeepSeek 236 10648 75.63 10292 73.10 356 3.34 Yes Original
5.2 Relationship Between Model Size and Overfit Detection
Figure 3 illustrates the scatter plot of the percentage performance difference versus the number of parameters, with a red dashed line representing the logarithmic fit ($\Delta_{1.0}=0.6318\cdot\ln\bigl{(}\text{\# Params}\bigr{)}+0.7920$). The significant log-linear relationship indicates that the performance difference increases with model size in a logarithmic fashion, suggesting diminishing returns as the number of parameters grows.
Figure 3: Scatter plot of the performance difference ($\Delta_{1.0}$) versus the number of model parameters (log scale). A logarithmic trendline is shown. Different colors represent different model families, highlighting how scaling affects robustness to perturbations.
Figure 4 plots the percentage performance difference ($\Delta_{1.0}$) for $\mu=1.0$ against the number of model parameters, with separate plots for models below and above 10B parameters (different x-axis scales). The data reveals a positive trend: larger models tend to exhibit greater performance degradation under textual perturbations. For example, models in the Gemma family show a progressive increase in $\Delta_{1.0}$ with higher parameter counts, while Llama models maintain low $\Delta_{1.0}$ values across scales. The dotted trend line further highlights this relationship.
Figure 4: Performance difference ($\Delta_{1.0}$) versus the number of model parameters for models with (a) fewer than and (b) more than 10 billion parameters. A trendline is shown, and different colors represent different model families, illustrating how model scale within each family relates to the impact of perturbations.
5.3 Relationship Between Model Accuracy and Overfit Detection
Figure 5 examines the relationship between baseline accuracy on the original prompts and the corresponding percentage difference in performance when evaluated on rephrased inputs. The plot clearly indicates that models with higher original accuracy tend to experience larger declines when exposed to prompt perturbations. For example, a model achieving over 80% accuracy on the original set shows one of the largest $\Delta_{1.0}$ values, while models with lower baseline accuracy exhibit only minor, often statistically insignificant, differences.
This observation highlights a paradox in current LLM evaluation: models that perform exceptionally well on standard benchmarks may be capitalizing on dataset-specific cues rather than demonstrating robust language understanding. The positive correlation between original accuracy and $\Delta_{1.0}$ underscores the need to carefully interpret high benchmark scores, as they might mask underlying vulnerabilities to prompt variations.
Figure 5: Scatter plot showing $\Delta_{1.0}$ for $\mu=1.0$ against the original accuracy of the model. Models within the same family are marked with the same color.
Table 3: Examples of how rephrasing affects LLM performance, illustrating potential overfitting to specific phrasing in the original MMLU dataset. The table shows original and rephrased questions, along with an explanation of why the model’s prediction changed. The examples are from Qwen2.5 (32B parameters).
Subject Original Question Rephrased Question Why the Model Was Wrong?
Professional Law “If the defendant is prosecuted for the man’s murder, he will most likely be found…” “If the defendant is charged with the man’s murder, what is the most probable outcome?” In legal contexts, terms like “prosecuted” and “found guilty/not guilty” are tied to specific legal standards. The rephrased question is more open-ended, leading the model to discuss outcomes like plea bargaining instead of focusing on the legal verdict.
Moral Disputes “Of the following social problems that could result from a genetic supermarket, which does Singer think is the least serious?” “Which of the following social issues arising from a genetic supermarket does Singer consider to be the least concerning?” The word “problems” was changed to “issues,” altering the model’s interpretation. “Issues” can broaden the context of "problems", causing the model to incorrectly interpret which concerns are least serious.
College Chemistry “Which of the following statements is not a reason why tetramethylsilane is used as a 1H chemical shift reference?” “Which of the following statements does not explain why tetramethylsilane is used as a reference for 1H chemical shifts?” The model may have overfit to the structure of the original question, particularly the phrase “is not a reason why,” as it directly signals the correct retrieval path. The rephrased version, with slight syntactic adjustments disrupts this memorization, leading to incorrect retrieval.
World Religions When did the first Jaina temples appear?.” At what point in time were the initial Jaina temples established?” The rephrased question shifts key terms (“When” to “At what point in time”), obscuring historical framing. The LLM fails to map this modified phrasing to the original temporal context.
These findings underscore the importance of evaluating LLMs under varied prompt formulations to ensure that improvements in benchmark performance reflect genuine advances in language understanding rather than overfitting.
6 Discussion
6.1 Why Do LLMs Overfit?
Table 3 highlights cases where LLMs answer the original questions correctly but fail on the rephrased versions. The failures suggest potential overfitting, where models overly rely on surface-level cues, memorized patterns, or specific terminologies. Overfitting in this context occurs because the model tends to associate certain question formats or keywords directly with answers instead of generalizing underlying concepts. Common root causes include shifts in terminology, subtle changes in phrasing that alter the semantic scope, and dependence on memorized patterns from training data.
6.2 Forget What You Know About LLM Evaluation
Ideally, LLMs should exhibit resilience when faced with variations in prompt wording and structure. In other words, robust LLMs are expected to maintain their performance regardless of how a question is phrased, thereby reflecting true language understanding rather than mere memorization. However, our experiments reveal a contrary trend: models that score highly on standard benchmarks often display heightened sensitivity to even minor alterations in prompt formulation. This behavior suggests that such models have implicitly overfitted to the specific linguistic patterns and structures characteristic of these benchmarks. As a result, when these surface-level cues are modified, performance declines, a phenomenon that underscores the paradox between high benchmark accuracy and genuine generalization.
Agnosticism to Benchmark Set.
Although we used MMLU as a demonstration, our approach is inherently dataset-agnostic. It can be applied to any benchmark by simply adapting the performance metric used to compare the original samples with their rephrased counterparts.
7 Conclusion
In this paper, we introduced a novel approach for detecting overfit to benchmarks datasets in LLMs by applying parametric transformations to these datasets. Our method revealed that many models rely heavily on surface features of public test sets, leading to significant performance drops when these features are altered. This finding underscores a critical insight: what appears to be robust performance may, in fact, be largely driven by memorization rather than true generalization.
We demonstrated the effectiveness of our approach across multiple LLM families. Notably, larger models tend to exhibit more pronounced performance declines under perturbation, while certain models (such as Llama) show greater stability. These observations suggest that training strategies and architectural choices play a significant role in mitigating overfitting, prompting a necessary rethinking of how we evaluate and benchmark LLMs.
By providing a practical, dataset-agnostic framework, our work equips the community with a powerful tool to uncover overfitting and to drive the development of benchmarks that better capture genuine generalization. Incorporating these parametric transformations into the evaluation process not only exposes hidden vulnerabilities in current LLMs but also suggests a way for the creation of more resilient models that can adapt to the evolving challenges of language tasks.
8 Limitations
While C-BOD serves as a promising framework for detecting overfitting in LLMs and has successfully identified overfitting in most evaluated models, it remains subject to several limitations. First, our approach primarily targets textual rephrasings that preserve semantic content. Consequently, it may overlook deeper forms of overfitting, such as factual inaccuracies or logical inconsistencies, which may require more specialized probing techniques. Moreover, incorporating $\mu$-based transformations into the training or fine-tuning loop can significantly increase computational cost. Iteratively rephrasing large datasets and retraining with multiple $\mu$ values imposes a heavy resource burden, which may not be feasible for LLMs or under restricted computational budgets. Future work should investigate more lightweight or partial-integration strategies. In summary, while C-BOD provides an effective means of detecting surface-level overfitting, further advancements are necessary to enhance its efficiency, scalability, and ability to capture more nuanced forms of model overfitting.
Acknowledgements
We used ChatGPT-4o for editing the language and refining the presentation of the text in this paper. The authors affirm that all research content and ideas are their own, and they take full responsibility for the final submitted manuscript.