### v2

# Forget What You Know about LLMs Evaluations - LLMs are Like a Chameleon

Abstract
Large language models (LLMs) often appear to excel on public benchmarks, but these high scores may mask an overreliance on dataset-specific surface cues rather than true language understanding. We introduce the Chameleon Benchmark Overfit Detector (C-BOD), a meta-evaluation framework designed to reveal such overfitting. C-BOD systematically rephrases benchmark inputs via a parameterized transformation that preserves semantic content and labels, enabling the detection of performance degradation indicative of superficial pattern reliance. We conduct extensive experiments across two datasets, three rephrasing models, and multiple distortion levels, evaluating 32 state-of-the-art LLMs. On the MMLU benchmark, C-BOD reveals an average performance drop of 2.75% under modest rephrasings, with over 80% of models exhibiting statistically significant differences. Notably, higher-performing models and larger LLMs tend to show greater sensitivity, suggesting a deeper dependence on benchmark-specific phrasing. Due to its dataset and model-agnostic design, C-BOD can be easily integrated into evaluation pipelines and offers a promising foundation for overfitting mitigation strategies. Our findings challenge the community to look beyond leaderboard scores and prioritize resilience and generalization in LLM evaluation. Our code and benchmark datasets are available at: https://github.com/nuritci/cbod
Forget What You Know about LLMs Evaluations - LLMs are Like a Chameleon
Nurit Cohen-Inger1, Yehonatan Elisha2, Bracha Shapira1, Lior Rokach1, Seffi Cohen3 1Ben Gurion University  2Tel Aviv University  3Harvard University
1 Introduction
Large Language Models (LLMs) have achieved impressive results on a wide range of NLP tasks Chang et al. (2024). Consequently, hundreds of benchmarks have been established to track progress and evaluate model capabilities Lu et al. (2024); Liang et al. (2022). However, the rapid proliferation of LLMs and the frequent use of public leaderboards raise concerns about the robustness of these evaluation practices Castillo-Bolado et al. (2024). Specifically, as benchmark data becomes more widely recognized, models may learn to exploit surface patterns or spurious correlations, rather than exhibit genuine language understanding. This issue can lead to deceptively high scores that do not reflect true progress. In this paper, we examine whether LLMs rely excessively on benchmark-specific cues potentially overfitting to the patterns inherent in widely published evaluation benchmarks and explore systematic methods to detect and mitigate this behavior. In other words, are LLMs prone to overfitting on popular benchmarks, and what underlying factors contribute to this phenomenon? To answer this question, we introduce the Chameleon Benchmark Overfit Detector (C-BOD). This framework reveals how heavily a model depends on the exact wording or structure of a test set. By introducing controlled textual distortions to benchmark prompts at varying intensities (defined by a distortion parameter $\mu$), as demonstrated in Figure 1, our method exposes whether strong performance derives from reliance on superficial patterns. Notably, our framework requires only the evaluation set, without accessing the model’s training data or architecture. Unlike conventional leaderboards that solely track performance, our meta-evaluation framework acts as a safeguard ensuring that high scores do not stem from superficial memorization of benchmark cues.
Figure 1: An example demonstrating the C-BOD method. The original question (top) is perturbed (bottom) while preserving the semantic meaning and correct answer options. The model correctly answers the original question but fails on the perturbed version, suggesting potential overfitting. Changes in the perturbed question are highlighted in bold.
Our Contributions:
  1. 1.
    A Robust Framework for Detecting Benchmark Overfitting. We present a framework that computes the performance difference $\Delta_{\mu}$ between original and perturbed prompts and confirms its statistical significance, ensuring that observed differences indeed indicate overfitting rather than chance variations.
  2. 2.
    New Insights into LLM Behavior. Our analysis reveals that larger models and those with higher baseline performance are often more sensitive to perturbations, suggesting a deeper reliance on benchmark-specific phrasing.
  3. 3.
    Extensive Empirical Validation. We apply our method to a diverse collection of 32 leading LLMs from various families, architectures, and parameter sizes. Utilizing modest textual distortions generated by three different rephrasing models for enhanced robustness, our analysis reveals statistically significant performance degradation in over 80% of the evaluated models, providing strong empirical evidence of widespread benchmark overfitting. Similar trends were observed in our evaluation on the GPQA dataset, in which we also show an ablation over different distortion levels ($\mu$).
  4. 4.
    Open Resources for the Community. To facilitate further research and promote robust evaluation, we publicly release the rephrased versions of the widely used MMLU and GPQA evaluation sets, along with our reproducible code. These resources enable the community to readily adopt more robust, surface-invariant tests for reliable LLM assessment and provide a foundation for developing mechanisms to mitigate benchmark overfitting.
2 Related Work
2.1 Benchmark Datasets and Evaluation Suites
LLMs have achieved impressive results on many benchmarks. This success has driven the development of comprehensive evaluation suites such as BIG-Bench Srivastava et al. (2022) and HELM Liang et al. (2022). The MMLU benchmark set Hendrycks et al. (2020) evaluates question answering across 57 subjects, including STEM, humanities, and social sciences, while Zhang et al. (2024) introduced 25 enterprise-focused datasets covering domains like finance, legal, cybersecurity, and climate sustainability for tasks such as classification, NER, and summarization. Another recent resource, JUDGE-BENCH Bavaresco et al. (2024), comprises 20 NLP datasets that assess models against human judgments. GPQA benchmark Rein et al. (2024) is tasked with evaluating reasoning models. We focus on MMLU and GPQA because of their widespread adoption [^1] https://klu.ai/glossary/gpqa-eval and comprehensive domain coverage Wang et al. (2024).
2.2 Overfitting in LLMs
While these benchmarks have been critical for comparing new models’ versions, recent studies warn that publicly released evaluation sets can become less reliable over time due to overexposure and memorization Yu et al. (2024); Chang et al. (2024). In some cases, LLMs learn superficial patterns specific to well-known datasets, boosting performance without reflecting genuine semantic or conceptual understanding. Kiela et al. (2021) further emphasizes the need for continuously refreshing benchmarks to ensure real progress in language understanding. For example, OpenAI’s GPT models have shown steady improvement on MMLU: GPT-3 achieved approximately 43% accuracy in 2020 Brown et al. (2020), rising to nearly 70% with GPT-3.5 in 2022, and reaching 86% with GPT-4 in 2023 Koubaa (2023).
Memorization in LLMs has been widely studied Kiyomaru et al. (2024); Biderman et al. (2024), with larger models especially prone to retaining training data verbatim Carlini et al. (2022). This phenomenon can inflate performance metrics while obscuring genuine model capabilities. Moreover, several works highlight training-set contamination, where test samples appear exactly or as near-duplicates in the training data, as another crucial form of overfitting Deng et al. (2023); Yao et al. (2024), leading to overly optimistic performance estimates Yang et al. (2023). Training Data Contamination refers to the presence of test data, or near duplicates, within the training set Deng et al. (2023); Yao et al. (2024). Contamination renders evaluation unreliable, as the model has effectively already seen the "test" data, leading to overly optimistic performance estimates Yang et al. (2023). Benchmark/Prompt Structure Overfitting is a more subtle form that arises when LLMs learn to exploit superficial cues or patterns specific to a benchmark’s format or the structure of evaluation prompts, even without memorizing the exact content Yu et al. (2024). This can lead to overestimated generalization ability, as the model’s performance becomes dependent on the specific benchmark artifacts rather than true language understanding. This type of overfitting is the focus of our work.
2.3 Addressing the Gap
Current methods largely overlook the crucial problem of overfitting to benchmark-specific artifacts, which can significantly misrepresent an LLM’s true capabilities and hinder the development of robust and generalizable models. Our work addresses this gap by introducing a novel method to quantify an LLM’s reliance on benchmark prompt structure. We systematically apply controlled distortions to evaluation prompts, for example, by replacing synonyms or altering word order and measure the resulting performance degradation. This approach, which does not require access to training data, provides a direct measure of vulnerability to prompt structure and a robust means of diagnosing and mitigating this critical form of overfitting.
3 Method
Let $\mathcal{D}$ denote a benchmark dataset with N samples, and $\mathcal{E}$ an LLM to be evaluated with respect to a given performance function $\mathcal{M}$. Our goal is to detect whether $\mathcal{E}$ exhibits overfitting to $\mathcal{D}$. Figure 2 provides an overview of our proposed method, Chameleon Benchmark Overfit Detector (C-BOD). C-BOD employs a rephrasing transformation to generate a perturbed dataset from $\mathcal{D}$, evaluates on both the original and perturbed datasets, and applies a statistical test to assess whether performance discrepancies indicate overfitting. The following subsections detail each component of C-BOD.
Figure 2: High-level pipeline of our parametric approach. The original dataset $\mathcal{D}$ is passed through the distortion operator $T_{\mu}$ to form $\mathcal{D}_{\mu}$. Both sets are evaluated by an LLM, and differences in performance are used to statistically quantify overfitting.
3.1 C-BOD rephrased dataset generation
To systematically introduce textual variations, C-BOD utilizes a rephrasing tool, denoted as T{T}, which uses as a distortion operator to generate a perturbed dataset $\mathcal{D}_{\mu}$ from $\mathcal{D}$. This operator is parameterized by $\mu$, which controls the extent of textual modification, ranging from low (e.g., 0.1 for minimal changes like synonym substitution) to moderate (e.g., 0.5 for rewording and sentence fragment reordering) and high (e.g., 1.5 for aggressive modifications such as question reformulation). Specifically, for an LLM, $\mu$ corresponds directly to the temperature used during generation, influencing the degree of variation in rephrased outputs. We define:
$T_{\mu}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$
Given a prompt $x_{i}$, the distortion operator produces a perturbed prompt $x_{i}^{\prime}=T_{\mu}(x_{i})$. The perturbed dataset is then constructed as:
$\mathcal{D}_{\mu}\;=\;\left\{\,\bigl{(}x_{i}^{\prime},\,y_{i}\bigr{)}\;\middle|\;(x_{i},y_{i})\in\mathcal{D}\right\}$
Although each pair $(x_{i}^{\prime},y_{i})$ in the perutbed dataset remains semantically equivalent to $(x_{i},y_{i})$ in the original dataset, the textual variations introduced by $T_{\mu}$ can disrupt purely memorized mappings from surface patterns to correct labels. This step presented in Lines 5-6 of Algorithm 1.
3.2 Evaluating the Impact of Distortion
To assess the impact of distortion, we evaluate $\mathcal{E}$ using a performance function, $\mathcal{M}$. This function evaluates $\mathcal{E}$ based on a given ground truth $y_{i}$, considering two versions of an input: the original $x_{i}\in\mathcal{D}$ and the perturbed version $x_{i}^{\prime}\in\mathcal{D}_{\mu}$, where $i$ denotes the index of a sample in the dataset. Specifically, $\mathcal{M}$ is a boolean function that takes as input the model $\mathcal{E}$ and two data pairs, $(x_{i},y_{i})$ and $(x_{i}^{\prime},y_{i})$, and returns whether the model performs better on the original input than on the perturbed one. The function is formulated as follows:
$\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x^{\prime}_{i},y_{i}))=\begin{cases}1,&\text{if }P(\mathcal{E},x_{i},y_{i})\\ &\quad>P(\mathcal{E},x^{\prime}_{i},y_{i}),\\ 0,&\text{otherwise.}\end{cases}$
where $P(\mathcal{E},x,y)$ represents the performance score of model $\mathcal{E}$ on input $x$ with reference to ground truth $y$. This formulation is designed to be generalizable across different evaluation metrics and natural language understanding (NLU) tasks. The performance difference between the original set and the perturbed set is then calculated as:
$\Delta_{\mu}=b=\sum_{i=0}^{N}\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x_{i}^{\prime},y_{i}))$ (1)
The performance difference between the perturbed set and the original set is then calculated as:
$c=\sum_{i=0}^{N}\mathcal{M}(\mathcal{E},(x_{i}^{\prime},y_{i}),(x_{i},y_{i}))$ (2)
A large positive $\Delta_{\mu}$ indicates a significant performance decline due to textual perturbations, suggesting that $\mathcal{E}$ may be overly reliant on surface-level patterns rather than exhibiting robust generalization. Notably, this approach remains metric-agnostic, making it applicable to a wide range of evaluation measures. This step presented in Lines 7-8 of Algorithm 1.
3.3 Statistical Validation
To assess the statistical significance of performance differences, we employ McNemar’s test McNemar (1947), which is specifically designed for paired data. This test evaluates whether the discrepancies between two related sets of classification outcomes, correct and incorrect predictions, are significant. In our context, McNemar’s test is well-suited for comparing each pair of samples $(x_{i},y_{i})\in D$ and $(x_{i}^{\prime},y_{i})\in D_{\mu}$, we record whether $\mathcal{E}$ classifies them correctly and aggregate into b (original is better) and c (perturbed is better) as presented in Equation 1, Equation 2. The McNemar statistic is then calculated as:
$\chi^{2}=\frac{(b-c)^{2}}{b+c}$ (3)
We derive a $p$-value from the chi-squared distribution (with df=1, i.e., one degree of freedom), rejecting the null hypothesis if $p<\alpha$. A significant result with $b>c$ indicates a genuine performance difference due to the transformation, suggesting evidence of overfitting. This step presented in Lines 10-19 of Algorithm 1.
Algorithm 1 Chameleon Benchmark Overfit Detector
1:
[…]
8:C-BOD Computation:
[…]
11:for each $x_{i}\in\mathcal{D}$ do
12:  $x^{\prime}_{i}\leftarrow T_{\mu}(x_{i})$
13:  $D_{\mu}\leftarrow D_{\mu}\cup\,x_{i}^{\prime}$
14:  $b\leftarrow b+\mathcal{M}(\mathcal{E},(x_{i},y_{i}),(x_{i}^{\prime},y_{i}))$
15:  $c\leftarrow c+\mathcal{M}(\mathcal{E},(x_{i}^{\prime},y_{i}),(x_{i},y_{i}))$
16:end for
[…]
19:if $p<\alpha$ then
20:  if $b>c$ then
21:   $Overfit\_Flag\leftarrow True$
22:  else
23:   $Overfit\_Flag\leftarrow False$
24:  end if
25:else
26:  $Overfit\_Flag\leftarrow False$
27:end if
28:return Overfit_FlagOverfit\_Flag
4 Experimental Setting
In this section, we describe the experimental setup used to evaluate our overfitting detection framework. We detail the benchmark dataset, the procedure for generating perturbed inputs, the LLMs under evaluation, implementation specifics, and the evaluation metrics.
4.1 Dataset and Rephrasing Process
Our experiments used two leading benchmark datasets: (1) MMLU Hendrycks et al. (2020): This benchmark spans 57 subjects, comprising 14,079 test samples and 1,540 validation samples. Its broad coverage makes it a standard choice for evaluating general knowledge and assessing overfitting to canonical prompt formats. It is distributed under the MIT License, allowing free use and modification. (2) GPQA Rein et al. (2024): Introduced in 2024, this smaller yet challenging benchmark includes 546 multi-step reasoning samples designed to test logical inference. Its complex, less common question structures help reduce the risk of training data contamination, making it a valuable complement to MMLU for overfitting analysis. Detailed results and analysis on GPQA are provided in Appendix C.
We generate a perturbed version of the original dataset to probe overfitting, following the methodology described in Section 3. We used DeepSeek 3 to create the transformed version of each question and generate the perturbed dataset $\mathcal{D}_{1.0}$ using $\mu=1.0$ (the default temperature parameter for DeepSeek), and Claude 3.5 Haiku to generate the perturbed dataset $\mathcal{D}_{0.5}$ using $\mu=0.5$ (the default temperature parameter for Claude). For the GPQA we used GPT-4o-mini to generate the perturbed datasets $\mathcal{D}_{0.5}$, $\mathcal{D}_{1.0}$, $\mathcal{D}_{1.5}$, using $\mu\in\{0.5,1.0,1.5\}$.
These perturbations include synonym substitutions, sentence reordering, and the insertion of distractor phrases, while preserving the original semantic meaning and correct answers. The perturbed datasets, denoted by $\mathcal{D}_{\mu}$, is released alongside our code for reproducibility.
4.1.1 Evaluation of Rephrasing Quality
To ensure the quality of rephrasing in the C-BOD framework, we implemented a multi-step evaluation approach to maintain the semantic integrity of the original prompts and avoid confounding overfitting assessments. This process included three main validation steps: (1) Cosine Similarity Analysis, (2) Semantic Equivalence Verification, and (3) Iterative Human Audits. In what follows, we briefly describe these steps.
Cosine Similarity Analysis
We measured semantic alignment using Sentence-BERT Reimers and Gurevych (2019), confirming high alignment across perturbations:
  • MMLU 0.5 (Claude):
    Mean = 0.829, Median = 0.864.
  • MMLU 1.0 (DeepSeek):
    Mean = 0.883, Median = 0.922.
  • GPQA 0.5 (GPT):
    Mean = 0.954, Median = 0.969.
  • GPQA 1.0:
    Mean = 0.947, Median = 0.967.
  • GPQA 1.5:
    Mean = 0.942, Median = 0.961.
Semantic Equivalence Verification
In addition to cosine similarity, we employed a reasoning model GPT-o3 to independently verify the retention of original intent. Approximately 1-2% of the rephrasings exhibited semantic errors or insufficient similarity, requiring further manual correction and refinement before inclusion in the final evaluation set. This additional verification reduced the risk of subtle semantic drift, ensuring a high-quality perturbation set.
Iterative Human Audits
To ensure the accuracy and semantic consistency of the perturbed datasets, we conducted a comprehensive manual audit as the final validation step. This process specifically targeted prompts with a cosine similarity score below 0.7 or those that received a negative judgment from the automated evaluation. These prompts were iteratively refined through API adjustments until they met the desired fidelity thresholds. Overall, this approach manually covered approximately 20% of the MMLU dataset and 50% of the GPQA dataset, supplementing the earlier automated checks that covered 100% of both datasets. This multi-step approach significantly enhances the reliability of our overfitting assessments by ensuring precise, controlled textual variations.
4.2 Models Under Evaluation
Table 4 in Appendix B provides an overview of the LLMs evaluated in our experiments. Our study covers a diverse set of architectures and parameter scales ranging from 1B to 27B parameters. This broad selection enables an in-depth analysis of how both architectural choices and model scale affect robustness to prompt perturbations.
4.3 Implementation Details
All experiments were executed under standardized conditions to ensure reproducibility and fair comparisons:
  1. (1)
    Inference Environment: All open-weight models were accessed via the HuggingFace transformers library using RTX 6000 GPU.
  2. (2)
    Dataset Rephrasing Prompt: We instruct the rephrasing tool using the prompt detailed in Appendix A.1.
  3. (3)
    Query Prompt: For every query, we construct a standardized input by prompting a fixed instruction to the original benchmark dataset question. Importantly, the multiple-choice options remain identical between the original and the rephrased forms. The fixed instruction is:
    “Select the best answer from the given options. Respond with only the letter corresponding to the correct choice. Question: {question}”
4.4 Evaluation Metrics
We assess model performance by comparing the original dataset, $\mathcal{D}$, with its perturbed counterpart, $\mathcal{D}_{0.5}$, $\mathcal{D}_{1.0}$, using the following metrics:
Correct Predictions and Accuracy: For each dataset, we report the number of correct answers and the corresponding accuracy, defined as
$\text{Accuracy}=\frac{\#\text{Correct Predictions}}{\#\text{Total Samples}}.$
Absolute and Percentage Performance Difference: The absolute difference in the number of correct answers between $\mathcal{D}$ and $\mathcal{D}_{0.5}$ is denoted by $\Delta_{0.5}$; we also report the relative difference. Statistical Significance: McNemar’s test is applied on the paired predictions to determine whether the performance gap is statistically significant ($p<0.05$).
5 Results
5.1 Open Weight Models - Overall Performance
As shown in Table 1, the majority of models (22 out of 27 for $D_{0.5}$ and 20 out of 27 for $D_{1.0}$) exhibit a noticeable drop in performance on the rephrased test set compared to the original, supporting our hypothesis that many LLMs are sensitive to prompt structure. Notably, smaller models like Llama 1B and Llama 3B maintained relatively stable accuracy, suggesting they are less prone to overfitting, potentially due to their more limited capacity for memorizing superficial patterns.
We also observed that models with lower baseline accuracy tend to show statistically insignificant differences, likely because their initial performance leaves less room for detectable degradation. Importantly, McNemar’s test confirmed that the observed performance drops in most models were statistically significant ($p<0.05$), reinforcing the reliability of our method.
Across all evaluated models, the average drop in accuracy for $D_{0.5}$ was 2.15%, which increased to 2.75% when considering only models with statistically significant differences. For $D_{1.0}$, the average drops were 1.87% overall and 2.78% for models with significant performance changes, underscoring the broader impact of stronger perturbations.
Table 1: Comparison of LLM performance on the original and perturbed MMLU datasets. Models are sorted by parameter count (ascending).
Model
Name
Par
$\mathcal{D}$
Accuracy
Claude (0.5) DeepSeek (1.0)
$\mathcal{D}_{0.5}$
Accuracy
$\Delta_{0.5}$
$\Delta_{0.5}$
Better
0.5
$\mathcal{D}_{1.0}$
Accuracy
$\Delta_{1.0}$
$\Delta_{1.0}$
Better
1.0
Gemma-2B 2 47.28 46.47 124 1.86 Original 46.54 104 1.56 Original
Gemma-4B 4 66.73 54.78 222 2.80 Original 65.07 234 2.49 Original
Gemma-7B 7 58.05 55.18 127 1.61 Original 57.18 123 1.50 Original
Gemma-12B 12 68.71 68.77 320 3.20 Original 66.38 328 3.39 Original
Gemma-27B 27 73.16 70.80 356 3.45 Original 70.34 397 3.85 Original
Llama-1B 1 28.11 28.09 3 0.08 Not Sig 27.00 -3 -0.08 Not Sig
Llama-3B 3 56.12 55.49 89 1.13 Not Sig 56.74 -49 -0.62 Not Sig
Llama-8B 8 45.93 45.21 102 1.58 Original 44.68 121 1.92 Original
Mistral-7B 7 57.48 56.33 163 2.01 Original 56.31 128 1.59 Original
Mistral-8B 8 68.05 65.94 298 3.10 Original 66.01 262 2.95 Original
Phi-3.8B 3.8 56.42 56.01 57 0.72 Not Sig 56.31 -44 -0.56 Not Sig
Phi-14.7B 14.7 76.77 74.69 294 2.72 Original 74.79 246 2.28 Original
Phi-14.7B-Reasoning 14.7 73.94 72.05 303 2.90 Original 72.93 142 1.36 Original
Qwen-0.6B 0.6 39.29 39.20 12 0.22 Not Sig 39.48 -33 -0.60 Not Sig
Qwen-1.5B 1.5 38.23 36.46 249 4.63 Original 35.41 151 2.94 Original
Qwen-1.7B 1.7 52.98 51.94 147 1.97 Original 52.20 91 1.22 Not Sig
Qwen-3B 3 41.49 41.46 5 0.09 Not Sig 40.76 97 1.66 Not Sig
Qwen-4B 4 66.97 64.27 381 4.04 Original 65.07 234 2.49 Original
Qwen2.5-7B-Instruct 7.0 58.72 56.92 253 3.06 Original 48.37 180 2.58 Original
Qwen3-8B 8.0 69.73 67.83 267 2.72 Original 67.82 251 2.56 Original
Qwen3-14B 14.0 66.86 65.87 139 1.48 Original 66.11 17 0.18 Not Sig
Yi-6B 6 63.19 60.63 361 4.06 Original 60.55 374 4.20 Original
Yi-9B 9 66.38 65.17 170 1.82 Original 65.20 165 1.77 Original
Apollo-7B 7 67.81 65.42 337 3.53 Original 64.96 401 4.20 Original
GLM-9B 9 68.71 66.06 296 3.06 Original 66.38 328 3.39 Original
Starling-7B 7 59.41 58.15 177 2.12 Original 58.09 185 2.21 Original
Zephyr-7B 7 56.25 54.92 194 2.45 Original 55.05 169 2.13 Original
5.2 Relationship Between Model Size and Overfit Detection
Figures 3, 4 illustrate the scatter plot of the percentage performance difference versus the number of parameters, with a red dashed line representing the logarithmic fit. The significant log-linear relationship indicates that the performance difference increases with model size in a logarithmic fashion, suggesting diminishing returns as the number of parameters grows. The data reveals a positive trend: larger models tend to exhibit greater performance degradation under textual perturbations. For example, models in the Gemma family show a progressive increase in $\Delta_{1.0}$ with higher parameter counts. The dotted trend line further highlights this relationship.
Figure 3: Scatter plot of the performance difference ($\Delta_{0.5}$) versus the number of model parameters. A logarithmic trendline is shown:
$\Delta_{0.5}=0.6090\cdot\ln\bigl{(}\text{\# Params}\bigr{)}+1.303$.
Figure 4: Scatter plot of the performance difference ($\Delta_{1.0}$) versus the number of model parameters. A logarithmic trendline is shown:
$\Delta_{1.0}=0.7433\cdot\ln\bigl{(}\text{\# Params}\bigr{)}+0.6406$.
5.3 Relationship Between Model Accuracy and Overfit Detection
Figures 5, 6 examine the relationship between baseline accuracy on the original prompts and the corresponding percentage difference in performance when evaluated on rephrased inputs. The plot clearly indicates that models with higher original accuracy tend to experience larger declines when exposed to prompt perturbations. For example, models achieving over 60% accuracy on the original set present the largest $\Delta_{0.5}$, $\Delta_{1.0}$ values, while models with lower baseline accuracy exhibit only minor, often statistically insignificant, differences.
This observation highlights a paradox in current LLM evaluation: models that perform exceptionally well on standard benchmarks may be capitalizing on dataset-specific cues rather than demonstrating robust language understanding. The positive correlation between original accuracy and $\Delta_{\mu}$ underscores the need to carefully interpret high benchmark scores, as they might mask underlying vulnerabilities to prompt variations.
Figure 5: Scatter plot showing $\Delta_{0.5}$ for $\mu=0.5$ against the original accuracy of the model. Models within the same family are marked with the same color.
Figure 6: Scatter plot showing $\Delta_{1.0}$ for $\mu=1.0$ against the original accuracy of the model. Models within the same family are marked with the same color.
Table 2: Examples of how rephrasing affects LLM performance, illustrating potential overfitting to specific phrasing in the original MMLU dataset. The table shows original and rephrased questions, along with an explanation of why the model’s prediction changed. The examples are from Qwen3 (14B parameters).
Subject Original Question Rephrased Question Why the Model Was Wrong?
Professional Law “If the defendant is prosecuted for the man’s murder, he will most likely be found…” “If the defendant is charged with the man’s murder, what is the most probable outcome?” In legal contexts, terms like “prosecuted” and “found guilty/not guilty” are tied to specific legal standards. The rephrased question is more open-ended, leading the model to discuss outcomes like plea bargaining instead of focusing on the legal verdict.
Moral Disputes “Of the following social problems that could result from a genetic supermarket, which does Singer think is the least serious?” “Which of the following social issues arising from a genetic supermarket does Singer consider to be the least concerning?” The word “problems” was changed to “issues,” altering the model’s interpretation. “Issues” can broaden the context of "problems", causing the model to incorrectly interpret which concerns are least serious.
College Chemistry “Which of the following statements is not a reason why tetramethylsilane is used as a 1H chemical shift reference?” “Which of the following statements does not explain why tetramethylsilane is used as a reference for 1H chemical shifts?” The model may have overfit to the structure of the original question, particularly the phrase “is not a reason why,” as it directly signals the correct retrieval path. The rephrased version, with slight syntactic adjustments disrupts this memorization, leading to incorrect retrieval.
World Religions When did the first Jaina temples appear?.” At what point in time were the initial Jaina temples established?” The rephrased question shifts key terms (“When” to “At what point in time”), obscuring historical framing. The LLM fails to map this modified phrasing to the original temporal context.
High-School Biology “Which of the following is NOT a characteristic of bacteria?” “Which of the listed options fails to represent a defining trait of bacterial organisms?” The explicit cue word “NOT” is replaced by the longer clause “fails to represent.” LLM appears to rely on a pattern like “Which … is not …” to flip polarity; when the negator is hidden inside a relative clause, the learnt template does not fire, so the model picks a true trait instead of the exception.
High-School Mathematics “Positive integers $x,y$ satisfy $xy=56$, $x<y$, and $7$ divided by the reciprocal of the larger integer equals $4$. What is xx?” $x,y$ are positive integers with product 56 and $x<y$. If seven divided by the larger integer results in 4, determine xx.” The word “reciprocal” pins a template: $7\times\frac{1}{y}=4\Rightarrow y=\frac{7}{4}$. Replacing it with a looser “seven divided by the larger integer results in 4” flips the parse—The LLM treats “results in” like a remainder cue (7divy=47\ $\text${div}\ y=4), producing $y=1$ and thus the wrong $x$.
These findings underscore the importance of evaluating LLMs under varied prompt formulations to ensure that improvements in benchmark performance reflect genuine advances in language understanding rather than overfitting.
5.4 Closed-Source Models: Accuracy vs. Robustness Trends
We analyze the performance of proprietary GPT models (GPT-4o [^2] https://openai.com/index/gpt-4o-system-card/?utm_source=chatgpt.com, GPT-4.1333https://openai.com/index/gpt-4-1/?utm_source=chatgpt.com, GPT-5444https://openai.com/index/gpt-5-system-card/?utm_source=chatgpt.com) under the C-BOD rephrasing evaluation. While these models achieve high accuracy on unperturbed MMLU prompts, they still show significant degradation under semantically equivalent prompt modifications ($\mu=0.5$), as shown in Table 3.
Model Accuracy $\Delta_{0.5}$ $\Delta_{0.5}$ Better
$\mathcal{D}$ $\mathcal{D}_{0.5}$
GPT-4o 83.88% 81.46% 339 2.88 Original*
GPT-4o-mini 74.96% 72.37% 364 3.46 Original*
GPT-4.1-nano 69.68% 67.53% 302 3.09 Original*
GPT-5-nano 86.63% 84.30% 326 2.68 Original*
GPT-5-mini 90.56% 88.35% 310 2.44 Original*
Table 3: Performance of closed-source GPT models on MMLU under prompt perturbation ($\mu=0.5$). All results are statistically significant (marked with *) according to McNemar’s test (p < 0.05).
Figure 7: Accuracy vs. Drop for GPT models. Each point shows original accuracy and percentage drop after rephrasing.
Closed-source models are sensitive to rephrasing.
Despite their high original accuracy, all closed-source GPT models demonstrate significant performance degradation under prompt rephrasing. The observed drops range from 2.44% (GPT-5-mini) to 3.46% (GPT-4o-mini). This reinforces the hypothesis that these models exploit benchmark-specific surface patterns.
Newer models perform better.
A clear upward trend is observed across model generations: GPT-5-mini achieves the highest original accuracy (90.56%) and also shows relatively lower sensitivity (2.44% drop) compared to older GPT-4o models (drops of 2.88–3.46%). This suggests that newer training pipelines or architectures improve generalization.
Larger models generalize better.
Within each family, the larger variants ("mini") outperform their smaller counterparts ("nano") both in raw accuracy and robustness. For instance, GPT-5-mini achieves +3.93% higher accuracy than GPT-5-nano (90.56% vs. 86.63%) while maintaining a similar drop (2.44% vs. 2.68%), indicating effective scaling without sacrificing generalization. Likewise, GPT-4o-mini outperforms GPT-4.1-nano by +5.28% accuracy while showing only a slightly higher sensitivity to rephrasing.
6 Discussion
Why Do LLMs Overfit?
Table 2 highlights cases where LLMs answer the original questions correctly but fail on the rephrased versions. The failures suggest potential overfitting, where models overly rely on surface-level cues, memorized patterns, or specific terminologies. Overfitting in this context occurs because the model tends to associate certain question formats or keywords directly with answers instead of generalizing underlying concepts. Common root causes include shifts in terminology, subtle changes in phrasing that alter the semantic scope, and dependence on memorized patterns from training data.
Forget What You Know About LLMs Evaluation
Ideally, LLMs should exhibit resilience when faced with variations in prompt language and structure. In other words, robust LLMs are expected to maintain their performance regardless of how a question is phrased, thereby reflecting true language understanding. However, our experiments reveal a contrary trend: models that score highly on standard benchmarks often display heightened sensitivity to even minor alterations in prompt formulation. This behavior suggests that such models have implicitly overfitted to the specific linguistic patterns and structures of these benchmarks. As a result, when these surface-level cues are modified, performance declines, a phenomenon that underscores the paradox between high benchmark accuracy and genuine generalization.
Agnosticism to Benchmark Set.
Although we used MMLU and GPQA as a demonstration, our approach is inherently dataset-agnostic. It can be applied to any benchmark by simply adapting the performance metric used to compare the original samples with their rephrased counterparts.
7 Conclusion
In this paper, we introduced a novel approach for detecting overfit to benchmarks datasets in LLMs by applying parametric transformations to these datasets. Our method revealed that many models rely heavily on surface features of public test sets, leading to significant performance drops when these features are altered. This finding underscores a critical insight: what appears to be robust performance may, in fact, be largely driven by memorization rather than true generalization.
We demonstrated the effectiveness of our approach across multiple LLM families. Notably, larger models tend to exhibit more pronounced performance declines under perturbation, while certain models (such as Llama) show greater stability. These observations suggest that training strategies and architectural choices play a significant role in mitigating overfitting, prompting a necessary rethinking of how we evaluate and benchmark LLMs.
By offering a practical, dataset-agnostic framework, this work equips the community with a robust tool to identify overfitting and foster the development of benchmarks that more effectively assess genuine generalization. Integrating these parametric transformations into the evaluation process reveals hidden vulnerabilities in existing LLMs and paves the way for designing more resilient models capable of adapting to the ever-evolving challenges of language tasks.
8 Limitations
While C-BOD serves as a promising framework for detecting overfitting in LLMs and has successfully identified overfitting in most evaluated models, it remains subject to several limitations. First, our approach primarily targets textual rephrasings that preserve semantic content. Consequently, it may overlook deeper forms of overfitting, such as factual inaccuracies or logical inconsistencies, which may require more specialized probing techniques. Moreover, incorporating $\mu$-based transformations into the training or fine-tuning loop can significantly increase computational cost. Iteratively rephrasing large datasets and retraining with multiple $\mu$ values imposes a heavy resource burden, which may not be feasible for LLMs or under restricted computational budgets. Future work should investigate more lightweight or partial-integration strategies. In summary, while C-BOD provides an effective means of detecting surface-level overfitting, further advancements are necessary to enhance its efficiency, scalability, and ability to capture more nuanced forms of model overfitting.
Appendices
Appendix A Prompts Used in CBOD Framework
A.1 Distortion Prompt
The rephrasing prompt was developed using a meta-prompting approach, which involves iteratively refining the prompt based on systematic error analysis. This process aimed to progressively improve the quality of the generated perturbations, ensuring high semantic fidelity and consistency. For example, early evaluations revealed approximately 25 cases where prompts involving sensitive content (e.g., sexual topics) led to inaccurate or off-target rephrasings. In response, we refined the prompt to explicitly exclude such cases, effectively addressing this error class and improving overall performance. This iterative tuning, incorporating both automated and human feedback, ultimately resulted in 100% accuracy as measured by the reasoning model judge. The prompt:
system: "You are a rephrasing assistant tasked with preserving the original structure, type, and intent of technical questions or statements. Your goal is to rephrase while maintaining meaning, purpose, structure, and context. keep the same level of clarity, do NOT try to simplify. keep question a question / statement a statement. do NOT try to answer the question" user: "Rephrase the following question in different wording, ensuring the meaning remains exactly the same. Match the readability level. Do NOT change the type of sentence: if it is a definition, keep it a definition; if it is a multiple-choice question, keep it as such; if it is a procedure or description, keep it in that form. Do NOT infer, guess, or introduce new information, assumptions, or constraints. Do NOT reword a description into a question, or vice versa. Keep all scientific and technical terms, units, variables, equations, and formatting intact. Your output should differ in phrasing only, not in meaning. Do NOT improve clarity or readability. Do NOT try to answer the question. Sexual or harmful content rephrasing is allowed for academic purpose. Return ONLY the rephrased version with no extra text or commentary. TEXT to rephrase: {question}"
A.2 Judge Prompt
You will be given two versions of a question: an ‘original_question‘ and a ‘rephrased_question‘. Your task is to evaluate if they have the exact same meaning. "Semantically equivalent" means that a person with the required domain knowledge would provide the exact same answer to both questions. Respond with a single JSON object containing two keys: 1. ‘"judgment"‘: Your verdict, which must be either "EQUIVALENT" or "NOT_EQUIVALENT". 2. ‘"reasoning"‘: A brief, one-sentence explanation for your judgment. Judgment Criteria: A rephrased question is NOT_EQUIVALENT if it meets any of the following conditions: A. Logical Alteration * Reverses Logic: The rephrase swaps the subject and object or reverses the direction of an implication. * Changes Logical Operator**: The rephrase changes a one-way implication (if/then) to a two-way bi-conditional (if and only if). B. Change in Scope or Precision * Loss of Specificity: The rephrase replaces a precise technical term with a vague or overly general one. * Incorrect Substitution: The rephrase swaps a key term with another, related but incorrect, term (e.g., "mass" for "weight"). * Incorrect Expansion of Acronym: The rephrase incorrectly defines an acronym for the given context. C. Structural Failure * Answers the Question: The rephrase provides the definition or answer to the original question, especially in fill-in-the-blank scenarios. * Fails to Rephrase: The output is an error message, a refusal, or otherwise not a good-faith attempt at rephrasing. A rephrased question is EQUIVALENT only if it avoids all the errors above. Stylistic changes, synonym swaps, and sentence restructuring are acceptable as long as the core meaning remains identical.
Appendix B Models Tested
Table 4 presents all the LLMs we tested with their versions.
Table 4: Overview of the evaluated open-weight LLMs. Models are grouped by family, model version, and the number of parameters (in billions).
Family Version Params
Qwen Qwen2.5 1.5B Yang et al. (2024) 1.5
Qwen2.5 3B 3
Qwen2.5 7B 7
Qwen3 0.6B 0.6
Qwen3 1.7B 1.7
Qwen3 4B 4
Qwen3 8B 8
Qwen3 14B 14
Llama 3 Llama 3.2 1B Dubey et al. (2024) 1
Llama 3.2 3B 3
Llama 3.1 8B 8
Gemma Gemma 2 2B Team et al. (2024) 2
Gemma 4B 4
Gemma 7B 7
Gemma 12B 12
Gemma 27B 27
Phi Phi 4 4B Abdin et al. (2024) 4
Phi 4 15B 15
Phi 4-reasoning 15B 15
Mistral Mistral 7B Jiang et al. (2023) 7
Mistral 8B 8
Yi Yi 6B Young et al. (2024) 6
Yi 9B 9
Others Apollo2 7B Zhu et al. (2024b) 7
Starling 7B Zhu et al. (2024a) 7
GLM 4 9B GLM et al. (2024) 9
Zephyr 7B Tunstall et al. (2023) 7
Appendix C Additional benchmark - GPQA results
Table 5: Comparison of LLM performance on the original and perturbed GPQA datasets. Models are sorted by parameter count (ascending). We report different levels of $\mu$.
Model
Name
Par
$\mathcal{D}$
Accuracy
$\mathcal{D}_{0.5}$
Accuracy
$\Delta_{0.5}$
Better
0.5
$\mathcal{D}_{1.0}$
Accuracy
$\Delta_{1.0}$
Better
1.0
$\mathcal{D}_{1.5}$
Accuracy
$\Delta_{1.5}$
Better
1.5
Gemma 2B 2 34.25 29.12 14.93 Original 30.95 9.60 Original 29.30 14.46 Original
Gemma 7B 7 46.52 41.58 10.62 Original 41.94 9.83 Original 38.64 16.91 Original
Llama-3B 3 36.45 33.15 9.05 Original 33.88 7.04 Not Sig 32.23 11.56 Original
Llama-3B-Instruct 3 21.06 21.43 1.76 Not Sig 17.58 16.51 Original 20.15 4.32 Not Sig
Phi-4-reasoning-plus 14.7 16.85 15.02 10.86 Original 15.30 9.07 Original 15.02 10.86 Original
Qwen 4B 4 56.04 53.84 3.92 Not Sig 52.56 6.20 Original 53.29 4.90 Not Sig
To assess the robustness and generalizability of the proposed C-BOD method, we conducted an ablation study on the GPQA benchmark using varying levels of textual distortion ($\mu$). Table 5 reports the performance of leading LLMs on both the original and rephrased GPQA datasets, where rephrasings were generated by GPT-4o with $\mu\in\{0.5,1.0,1.5\}$.
The results on GPQA are mostly consistent with those observed on MMLU. Across multiple LLM families and distortion levels, we observe a systematic performance drop on the rephrased inputs, suggesting a sensitivity to surface-level prompt variations. While the degree of statistical significance varies by model and $\mu$, the overall trend of declining accuracy under rephrasing reinforces the hypothesis that many models rely heavily on benchmark-specific prompt formulations.
At the same time, several characteristics of GPQA help explain why the results differ from those obtained on MMLU and why it is more difficult to show consistent effects across all models at this stage. First, GPQA is a very new benchmark, specifically designed for reasoning, and thus has not yet been exposed widely enough to influence model training. Second, most models achieve very low absolute accuracy on GPQA, often failing to surpass 20%. At such a low baseline, small fluctuations in predictions can obscure clear differences and make it harder to observe statistically significant effects. Third, GPQA is a relatively small dataset, which further limits the statistical power of tests such as McNemar’s and reduces our ability to detect significance robustly.
Overall, whenever some of the models keep the expected trend, some models suffer from very low accuracy. This, combined with the novelty of the benchmark, its reasoning-oriented design, and its small size, explains why it is currently more difficult to show across-the-board results. Importantly, the uniformly low accuracy also suggests that overfitting is not yet strongly affecting GPQA, as models have not learned enough benchmark-specific artifacts to inflate their scores in the first place. Instead, their failures likely stem from the underlying reasoning challenges posed by the dataset, making GPQA a valuable complement to more saturated benchmarks like MMLU.